OpenStudy (anonymous):

Part 1: Solve the following system of equations. Please show your work. x - y = 16 2x + y = 2 Part 2: Explain which method you chose to solve the system and why you felt it was the best choice.

OpenStudy (hesan):

X=6, Y=-10

OpenStudy (hesan):

Part 2: System of two equations involving two variables..

OpenStudy (anonymous):

firstly add the two equations ,you get 3x=18 x=6 and put the value of x in first equation to get the value of y 6-y=16 -y=16-6 -y=10 y=-10